A robot uses a special phase difference-based laser sensor to determine the distance from obstacles in his way. This sensor continuously outputs the distance but due to a glitch in the sensor electronics, it randomly returns an incorrect value at an average rate of 3 times per second. What is the likelihood that the sensor does not return any incorrect value within the first minute of operation?


1
Expert's answer
2021-09-20T01:45:14-0400

The probability of the sensor returning any incorrect value in a second "=\frac{3}{60}"

The probability of the sensor returning any correct value in a second "=\frac{57}{60}"

The likelihood that the sensor does not return any incorrect value within the first minute of operation

"=\frac{57}{60}=\frac{19}{20}"